OpenStudy (anonymous):

A runner starts from rest and speeds up with constant acceleration. If she has gone a distance of 30 m at the point when she reaches a speed of 8 m/s, what is her acceleration?

OpenStudy (kropot72):

First we need to find the time it takes to accelerate to a speed of 8 m/s. The average speed is \[\frac{0+8}{2}=4\ m/s\] The time taken to travel 30 m at an average speed of 4 m/s is \[\frac{30}{4}=7.5\ seconds\] The acceleration is therefore: \[a=\frac{8}{7.5}=\ (you\ can\ calculate)\ m/s ^{2}\]